OpenStudy (o-b-server):

Can someone help me with a graphing and functions question? It's just one question! Medal for best answer!

OpenStudy (o-b-server):

Find the third, sixth, and eighth terms of the sequence A(n) = 5 + (n – 1)(–2) A. 1, –5, –9 B. 1, –5, –11 C. 3, –3, –9 D. 1, –3, –9 I'm not very good with terms of a sequence so if you could also help explain it to me that'd be great!

OpenStudy (animeasgardian):

Ok, I'm not really good at this either, but I'll try to help. Basically what I do is substitute n for one of the answer choices.

OpenStudy (o-b-server):

Substitute? I don't get how'd that work.

OpenStudy (animeasgardian):

Ok....ignore everything i just said

OpenStudy (animeasgardian):

What you do is instead of n, it's 3, 6, and 8

OpenStudy (animeasgardian):

@O-B-Server

OpenStudy (o-b-server):

Okay? Well I get that I need to find the 3rd, 6th, and 8th term but it still doesn't make sense.

OpenStudy (animeasgardian):

Solve this: A(3)=5+(3-1)(-2)

OpenStudy (o-b-server):

Hm... Would the answer be 3?

OpenStudy (animeasgardian):

Nope.

OpenStudy (animeasgardian):

First you would subtract 3-1, which is 2

OpenStudy (animeasgardian):

now you have a(5)=5+2(-2)

OpenStudy (animeasgardian):

solve that

OpenStudy (o-b-server):

Wait how does the number next to A become a 5?

OpenStudy (animeasgardian):

Oh i'm sorry that was a typo, it's 3

OpenStudy (o-b-server):

Okay so 5+2=7 so we'd multiply it with -2 right?

OpenStudy (animeasgardian):

its a(3)=5+2(-2)

OpenStudy (animeasgardian):

You would do 2 * -2

OpenStudy (animeasgardian):

Then add 5

OpenStudy (o-b-server):

But that's what i did. 2 * -2 equals -4 and adding that to five equals 1

OpenStudy (animeasgardian):

Yes that's correct

OpenStudy (o-b-server):

Then what does the 3 next to the A for?

OpenStudy (animeasgardian):

Now we can eliminate C

OpenStudy (animeasgardian):

Oh 3 next to the a....I'm not sure what it means.

OpenStudy (o-b-server):

Ah it's fine. Thanks though.

OpenStudy (animeasgardian):

Ok now next do a(6)=5+(6-1)(-2)

OpenStudy (animeasgardian):

Are you still there @O-B-Server